# Online Estimation with Rolling Validation: Adaptive Nonparametric Estimation with Streaming Data

### v3

@@ Appendix F Technical Details for Choosing Weighting Exponents
This section offers more technical details for the discussion in Section 3.1.
@@
Consider two sequences $\alpha_{i}=Ai^{-a}$ and $\beta_{i}=Bi^{-b}$. Suppose $0<A<B$ and $10\leq a<b<$ 1. Then
$\beta_{i}\leq\alpha_{i}\text{ when }i\geq(B/A)^{(b-a)^{-1}}.$ (99)
Moreover, let $\xi>0$, we have
$\sum_{i=0}^{n}i^{\xi}\beta_{i}\leq\sum_{i=0}^{n}i^{\xi}\alpha_{i}\text{ when }n\geq(B/A)^{(b-a)^{-1}}M_{n},$ (100)
@@
$M_{n}=\left(\frac{\xi+1-a}{\xi+1-b}\right)^{1/(b-a)}\left(1+\frac{\xi-b+1}{n}\right)^{1/(b-a)}.$
@@
It is direct to verify equation (99). Now we show the details of deriving (100).
@@
We just need to solve the condition for
$\frac{A}{\xi-a+1}n^{\xi-a+1}\geq\frac{B}{\xi-b+1}(n+1)^{\xi-b+1}+Bn^{\xi-b}.$ (103)
It is equivalent to
$n\geq(B/A)^{(b-a)^{-1}}\left(\frac{\xi+1-a}{\xi+1-b}\right)^{1/(b-a)}\left(1+\frac{\xi-b+1}{n}\right)^{1/(b-a)}$ (104)
@@
The ratio between the thresholds Lemma F.2, $M_{n}$, is the delay ratio. We analyze the main component that does not vanish with increasing $n$. The solution of $\left(\frac{\xi+1-a}{\xi+1-b}\right)^{1/(b-a)}\leq T$ for some $T>1$ is
@@
For each fixed pair of $(a,b)$, we can treat $\xi$ as a function of the target delay ratio $T$. This function with some typical choices of the $a,b$-convergence rates is plotted in Figure F.1. We can observe that in order to achieve a smaller delay ratio $T$, we need to set the diverging weight exponent $\xi$ to be a larger value. Specifically, if we take $\xi=1$, we can control the ratio to be less than $2.4$; if it is set to be a larger number such as $2$ or $3$, the delay can be further shortened to $1.6$ or $1.4$.
@@
$\left(\frac{\xi+1+\delta-b}{\xi+1-b}\right)^{\delta^{-1}}$ (107)
@@
$\left(\frac{\xi+1+\delta-b}{\xi+1-b}\right)^{\delta^{-1}}\leq\exp\left(\xi^{-1}\right)$ (108)
@@ F.1 Efficiency loss at $I_{F}$
$E\left[r_{1,i}\right]/E\left[r_{2,i}\right]$ $=\frac{AI_{F}^{-a}}{BI_{F}^{-b}}=\frac{\xi+1-a}{\xi+1-b}$ (110)
@@ G.1 Regarding Batch Sieve Estimators
$E\{\hat{f}(X)-f(X)\}^{2}$ (111)
$=$ $E\left\{\sum_{j=1}^{J}\hat{\beta}_{j}\phi_{j}(X)-\sum_{j=1}^{J}\beta_{j}\phi_{j}(X)-\sum_{j=J+1}^{\infty}\beta_{j}\phi_{j}(X)\right\}^{2}$
$=$ $E\left\{\sum_{j=1}^{J}(\hat{\beta}_{j}-\beta_{j})\phi_{j}(X)-\sum_{j=J+1}^{\infty}\beta_{j}\phi_{j}(X)\right\}^{2}$
@@
$E\left(\hat{\beta}_{j}-\beta_{j}\right)^{2}=E\left(n^{-1}\sum_{i=1}^{n}Y_{i}\phi_{j}\left(X_{i}\right)-E\left[Y\phi_{j}(X)\right]\right)^{2}=n^{-1}\operatorname{Var}\left(Y\phi_{j}(X)\right).$
Under very mild conditions on the noise distribution and basis functions, each $E\left(\hat{\beta}_{j}-\beta_{j}\right)^{2}$ can be bounded by $n^{-1}$ times a constant. For the approximation error:
$\sum_{j=J+1}^{\infty}\beta_{j}^{2}\leq J^{-2s}\sum_{j=J+1}^{\infty}j^{2s}\beta_{k}^{2}=J^{-2s}Q^{2},$ (112)
@@
$\mathcal{K}(a,b)$ $=\sinh(1)^{-1}\cdot\cosh(a\wedge b))\cdot\cosh(1-a\vee b)$ (113)
@@
$\eta_{i}=\beta_{i}-\beta^{\prime}_{i}.$ (116)
@@
$\eta_{j}=\gamma Y_{i}X_{i}+\left(I-\gamma X_{i}X_{i}^{\top}\right)\beta_{i-1}-\gamma Y^{\prime}_{i}X^{\prime}_{i}+\left(I-\gamma X^{\prime}_{i}X_{i}^{\prime\top}\right)\beta_{i-1}$ (117)
@@
$\eta_{i}$ $=(I-\gamma\|X_{i}\|^{2}H_{i})\eta_{i-1}$ (118)
$=H_{i}^{\perp}\eta_{i-1}+(1-\gamma\|X_{i}\|^{2})H_{i}\eta_{i-1},$ (119)
@@
$\|\eta_{i}\|^{2}$ $=\|H_{i}^{\perp}\eta_{i-1}\|^{2}+(1-\gamma\|X_{i}\|^{2})^{2}\|H_{i}\eta_{i-1}\|^{2}$ (120)
$\stackrel{{\scriptstyle(I)}}{{\leq}}\|H_{i}^{\perp}\eta_{i-1}\|^{2}+(1-\gamma\|X_{i}\|^{2})\|H_{i}\eta_{i-1}\|^{2}$ (121)
$\leq\|\eta_{i-1}\|^{2}-\gamma\|X_{i}\|^{2}\|H_{i}\eta_{i-1}\|^{2}$ (122)
@@
$\frac{\|\eta_{i}\|^{2}}{\|\eta_{i-1}\|^{2}}\leq 1-\gamma\|X_{i}\|^{2}\frac{\|H_{i}\eta_{i-1}\|^{2}}{\|\eta_{i-1}\|^{2}}=1-\gamma\frac{\eta_{i-1}^{\top}X_{i}X_{i}^{\top}\eta_{i-1}}{\|\eta_{i-1}\|^{2}}$ (123)
@@
$E\left[\left.\frac{\|\eta_{i}\|^{2}}{\|\eta_{i-1}\|^{2}}\right|F^{i-1}\right]$ $\leq 1-\gamma\|\eta_{i-1}\|^{-2}\eta_{i-1}^{\top}E[X_{i}X_{i}^{\top}]\eta_{i-1}$ (124)
@@
In step $(II)$ we used $\lambda_{min}(E[X_{i}X_{i}^{\top}])\geq\underline{\lambda}>0$. Now we have established some “exponential contraction” properties for the sequence of vectors $\{\eta_{i}\}$. Then we can apply Lemma G.1 (with parameter $a$ therein equals to 00) to conclude that the averaged coefficient vector $\hat{\eta}_{i}=i^{-1}\sum_{k=1}^{i}\eta_{k}$ has magnitude:
$E[\|\hat{\eta}_{i}\|^{2}|F^{j}]\lesssim i^{-2}\|\eta_{j}\|^{2}.$ (125)
@@
$E\left[\left\{\nabla_{j}\hat{f}_{i-1}\left(X_{i}\right)\right\}^{2}\mid F^{j}\right]$ $=E\left[\left\{\hat{f}_{i-1}\left(X_{i};Z_{i}\right)-\hat{f}_{i-1}\left(X_{i};Z_{i}^{j}\right)\right\}^{2}\mid F^{j}\right]$ (126)
$=E\left[\left\{X_{i}^{\top}\hat{\beta}_{i-1}-X_{i}^{\top}\hat{\beta}_{i-1}^{\prime}\right\}^{2}\mid F^{j}\right]$
@@
$=E\left[\left.\{X_{i}^{\top}\hat{\eta}_{i-1}\}^{2}\right|F^{j}\right]$
$\leq E\left[\left.\|X_{i}\|^{2}\|\hat{\eta}_{i-1}\|^{2}\right|F^{j}\right],$
Under the assumption that $E[\|X_{i}\|^{2}]\leq R^{2}$,
$E\left[\left\{\nabla_{j}\hat{f}_{i-1}\left(X_{i}\right)\right\}^{2}\mid F^{j}\right]$ $\leq R^{2}E\left[\left\|\hat{\eta}_{i-1}\right\|^{2}\mid F^{j}\right]$ (127)
@@
$E\left[\left\{\nabla_{j}\hat{f}_{i-1}\left(X_{i}\right)\right\}^{2}\mid F^{j}\right]\lesssim i^{-2},$ (129)
@@
$\beta_{j}=\beta_{j-1}^{\wedge}+\gamma_{j}\left(Y_{j}-\bm{\phi}_{j}^{\top}\beta_{j-1}^{\wedge}\right)D_{j}\phi_{j},$ (130)
And we have a similar updating rule for $\beta_{j}^{\prime}$ using sample $\left(X_{j}^{\prime},Y_{j}^{\prime}\right)$ :
$\beta_{j}^{\prime}$ $=\beta_{j-1}^{\prime\wedge}+\gamma_{j}\left(Y_{j}-\bm{\phi}_{j}^{\prime\top}\beta_{j-1}^{\prime\wedge}\right)D_{j}\bm{\phi}_{j}^{\prime}$ (131)
@@
$\eta_{j}=\gamma_{j}\left(Y_{j}-\bm{\phi}_{j}^{\top}\beta_{j-1}^{\wedge}\right)D_{j}\bm{\phi}_{j}-\gamma_{j}\left(Y_{j}-\bm{\phi}_{j}^{\prime\top}\beta_{j-1}^{\prime\wedge}\right)D_{j}\bm{\phi}_{j}^{\prime}.$ (132)
@@
$\eta_{i}=\eta_{i-1}^{\wedge}-\gamma_{i}D_{i}\bm{\phi}_{i}\bm{\phi}_{i}^{\top}\eta_{i-1}^{\wedge}$ (133)
@@
$D_{i}^{-1/2}\eta_{i}=D_{i}^{-1/2}\eta_{i-1}^{\wedge}-\gamma_{i}D_{i}^{1/2}\bm{\phi}_{i}\left(D_{i}^{1/2}\bm{\phi}_{i}\right)^{\top}D_{i}^{-1/2}\eta_{i-1}^{\wedge}$ (134)
@@
$\|D_{j}^{1/2}\bm{\phi}_{j}\|^{2}=\sum_{k=1}^{J_{j}}(k^{-\omega}\phi_{k}(X_{j}))^{2}\leq C\sum_{k=1}^{\infty}k^{-2\omega}<\infty,$ (137)
@@
$\left\|\theta_{i}\right\|^{2}$ $=\left\|\theta_{i-1}^{\wedge}\right\|^{2}+\gamma_{i}^{2}\left(\theta_{i-1}^{\wedge}\right)^{\top}W_{i}W_{i}^{\top}W_{i}W_{i}^{\top}\theta_{i-1}^{\wedge}-2\gamma_{i}\left(\theta_{i-1}^{\wedge}\right)^{\top}W_{i}W_{i}^{\top}\theta_{i-1}^{\wedge}$ (138)
@@
In step $(II)$ above we used $\gamma_{i}\|W_{i}\|^{2}\leq 1$. Take conditional expectation
$E[\|\theta_{i}\|^{2}|F^{i-1}]$ $\leq\|\theta^{\wedge}_{i-1}\|^{2}-\gamma_{i}\theta_{i-1}^{{}^{\wedge}\top}E[W_{i}W_{i}^{\top}]\theta^{\wedge}_{i-1}$ (139)
$=\|\theta_{i-1}\|^{2}-\gamma_{i}\theta_{i-1}^{{}^{\wedge}\top}D_{i}^{1/2}E[\bm{\phi}_{i}\bm{\phi}_{i}^{\top}]D_{i}^{1/2}\theta^{\wedge}_{i-1}$
@@
$E[\|\theta_{i}\|^{2}/\|\theta_{i-1}\|^{2}|F^{i-1}]\leq 1-\gamma_{i}\frac{\theta_{i-1}^{{}^{\wedge}\top}D_{i}^{1/2}E[\bm{\phi}_{i}\bm{\phi}_{i}^{\top}]D_{i}^{1/2}\theta^{\wedge}_{i-1}}{\|\theta^{\wedge}_{i-1}\|^{2}}$ (140)
By our assumption, $\lambda_{min}(E[\bm{\phi}_{i}\bm{\phi}_{i}^{\top}])$ is strictly larger than a constant (we postpone the proof after equation (145)). Also note that $D_{i}^{1/2}={\rm Diag}(1^{-\omega},...,J_{i}^{-\omega})$ has eigenvalues greater than $J_{i}^{-\omega}=i^{-\zeta\omega}$. So overall we have
$\frac{\theta_{i-1}^{{}^{\wedge}\top}D_{i}^{1/2}E[\bm{\phi}_{i}\bm{\phi}_{i}^{\top}]D_{i}^{1/2}\theta^{\wedge}_{i-1}}{\|\theta^{\wedge}_{i-1}\|^{2}}\geq i^{-2\zeta\omega}.$ (141)
Then we have a exponential contraction recursion:
$E[\|\theta_{i}\|^{2}/\|\theta_{i-1}\|^{2}|F^{i-1}]\leq 1-\gamma_{i}i^{-2\zeta\omega}=1-i^{-2\zeta\omega-\zeta}\leq\exp(-i^{-\alpha}),$ (142)
@@
$E[\|\hat{\theta}_{i}\|^{2}|F^{j}]\lesssim i^{-2}j^{2\alpha}\|\theta_{j}\|^{2}.$ (143)
@@
$E[\|\hat{\theta}_{i}\|^{2}|F^{j}]\lesssim i^{-2}j^{2\alpha-2\zeta}=i^{-2}j^{4\omega\zeta}$ (144)

### v4

@@ Appendix F Technical Details for Choosing Weighting Exponents
This section offers more technical details for the discussion in Section 3.1. We will use the notation
$T(a,b,\xi)=\left(\frac{\xi+1-a}{\xi+1-b}\right)^{1/(b-a)}.$
@@
Consider two sequences $\alpha_{i}=Ai^{-a}$ and $\beta_{i}=Bi^{-b}$ for $i\geq 1$. Suppose $0<A<B$ and $10\leq a<b<$ 1. Then $\beta_{i}\leq\alpha_{i}$ when
$i\geq\left(\frac{B}{A}\right)^{1/(b-a)}.$
Moreover, for any $\xi\geq 0$, we have
$\sum_{i=1}^{n}i^{\xi}\beta_{i}\leq\sum_{i=1}^{n}i^{\xi}\alpha_{i}$
whenever $n$ satisfies
$n\geq\left(\frac{B}{A}\right)^{1/(b-a)}T(a,b,\xi)+r_{n},$
and
$\frac{B(\xi-a+1)}{An}+\frac{1}{n^{\xi-b+1}}\leq\frac{B}{A}T^{b-a}(a,b,\xi),$
@@
$r_{n}$ $=\frac{1}{b-a}\cdot 2^{1/(b-a)-1}\left(\frac{B}{A}\right)^{1/(b-a)-1}T^{1-(b-a)}\left(\frac{B(\xi-a+1)}{An}+\frac{1}{n^{\xi-b+1}}\right)$
$=O\left(\frac{1}{n}+\frac{1}{n^{\xi-b+1}}\right).$
@@
The first claim is straightforward to verify. We now detail the derivation of the second one. By assumption, $\xi-b>-1$. We will use the monotone–rectangle argument:
@@
So,
$\sum_{i=1}^{n}i^{\xi}\beta_{i}$ $=\beta_{1}+\sum_{i=2}^{n}Bi^{\xi-b}$
[…]
$=B\left(1+\frac{n^{\xi-b+1}-1}{\xi-b+1}\right)$
$\leq B\left(\frac{n^{\xi-b+1}}{\xi-b+1}+n^{\xi-b}\right).$
When $\xi-b>0$,
$\sum_{i=1}^{n}i^{\xi}\beta_{i}$ $=\sum_{i=1}^{n-1}i^{\xi}\beta_{i}+Bn^{\xi-b}$
[…]
$=B\left(\frac{n^{\xi-b+1}-1}{\xi-b+1}+n^{\xi-b}\right)$
$\leq B\left(\frac{n^{\xi-b+1}}{\xi-b+1}+n^{\xi-b}\right).$
Similarly, for $\alpha_{i}$, when $-1<\xi-a\leq 0$,
$\sum_{i=1}^{n}i^{\xi}\alpha_{i}$ $=\sum_{i=1}^{n-1}Ai^{\xi-a}+An^{\xi-a}$
[…]
$=A\left(\frac{n^{\xi-a+1}-1}{\xi-a+1}+n^{\xi-a}\right)$
$>A\frac{n^{\xi-a+1}-1}{\xi-a+1}.$
When $\xi-a>0$,
$\sum_{i=1}^{n}i^{\xi}\alpha_{i}$ $=A+\sum_{i=2}^{n}Ai^{\xi-a}$
[…]
$=A\left(1+\frac{n^{\xi-a+1}-1}{\xi-a+1}\right)$
$>A\frac{n^{\xi-a+1}-1}{\xi-a+1}.$
Hence, to ensure $\sum_{i=1}^{n}i^{\xi}\beta_{i}\leq\sum_{i=1}^{n}i^{\xi}\alpha_{i}$ we just need to find a sufficient condition for
$A\frac{n^{\xi-a+1}-1}{\xi-a+1}\geq B\left(\frac{n^{\xi-b+1}}{\xi-b+1}+n^{\xi-b}\right).$
Solve this inequality:
$n\geq\left(\frac{B}{A}\cdot\frac{\xi+1-a}{\xi+1-b}+\frac{B(\xi-a+1)}{An}+\frac{1}{n^{\xi-b+1}}\right)^{1/(b-a)}.$
We will use the inequality $(x+e)^{p}\leq x^{p}+p2^{p-1}x^{p-1}e$ for $x>0$, $0<e\leq x$, and $p>1$. Define
$e_{n}:=\frac{B(\xi-a+1)}{An}+\frac{1}{n^{\xi-b+1}}.$
When $e_{n}<\frac{B}{A}\cdot\frac{\xi+1-a}{\xi+1-b}$, a further sufficient condition is
$n\geq\left(\frac{B}{A}\cdot\frac{\xi+1-a}{\xi+1-b}\right)^{1/(b-a)}+r_{n}.$
@@
The ratio between the thresholds
$(B/A)^{1/(b-a)}$
and
$(B/A)^{1/(b-a)}T(a,b,\xi)+r_{n}$
in Lemma F.2 is essentially $T(a,b,\xi)$ up to an $O(r_{n})$ term. We will analyze this main component that does not vanish with an increasing $n$. The solution of $T(a,b,\xi)\leq t$ for some $t>1$ is
@@
For each fixed pair of $(a,b)$, we can treat $\xi$ as a function of the target delay ratio $t$. This function with some typical choices of the $a,b$-convergence rates is plotted in Figure F.1. We can observe that in order to achieve a smaller delay ratio $t$, we need to set the diverging weighting exponent $\xi$ to be a larger value. Specifically, if we take $\xi=1$, we can control the ratio to be less than $2.4$; if it is set to be a larger number such as $2$ or $3$, the delay can be further shortened to $1.6$ or $1.4$.
@@
$\left(\frac{\xi+1+\delta-b}{\xi+1-b}\right)^{\delta^{-1}}$
@@
$\left(\frac{\xi+1+\delta-b}{\xi+1-b}\right)^{\delta^{-1}}\leq\exp\left(\xi^{-1}\right)$
@@ F.1 Efficiency loss at $I_{F}$
$\mathbb{E}\left[r_{i,1}\right]/\mathbb{E}\left[r_{i,2}\right]$ $=\frac{AI_{F}^{-a}}{BI_{F}^{-b}}=\frac{\xi+1-a}{\xi+1-b}$
@@ G.1 Regarding Batch Sieve Estimators
$\mathbb{E}\{\hat{f}(X)-f(X)\}^{2}$
$=$ $\mathbb{E}\left\{\sum_{j=1}^{J}\hat{\beta}_{j}\phi_{j}(X)-\sum_{j=1}^{J}\beta_{j}\phi_{j}(X)-\sum_{j=J+1}^{\infty}\beta_{j}\phi_{j}(X)\right\}^{2}$
$=$ $\mathbb{E}\left\{\sum_{j=1}^{J}(\hat{\beta}_{j}-\beta_{j})\phi_{j}(X)-\sum_{j=J+1}^{\infty}\beta_{j}\phi_{j}(X)\right\}^{2}$
@@
$\mathbb{E}\left(\hat{\beta}_{j}-\beta_{j}\right)^{2}=\mathbb{E}\left(n^{-1}\sum_{i=1}^{n}Y_{i}\phi_{j}\left(X_{i}\right)-\mathbb{E}\left[Y\phi_{j}(X)\right]\right)^{2}=n^{-1}\operatorname{Var}\left(Y\phi_{j}(X)\right).$
Under very mild conditions on the noise distribution and basis functions, each $\mathbb{E}\left(\hat{\beta}_{j}-\beta_{j}\right)^{2}$ can be bounded by $n^{-1}$ times a constant. For the approximation error:
$\sum_{j=J+1}^{\infty}\beta_{j}^{2}\leq J^{-2s}\sum_{j=J+1}^{\infty}j^{2s}\beta_{k}^{2}=J^{-2s}Q^{2},$
@@
$\mathcal{K}(a,b)$ $=\sinh(1)^{-1}\cdot\cosh(a\wedge b))\cdot\cosh(1-a\vee b)$
@@
$\eta_{i}=\beta_{i}-\beta^{\prime}_{i}.$
@@
$\eta_{j}=\gamma Y_{i}X_{i}+\left(I-\gamma X_{i}X_{i}^{\top}\right)\beta_{i-1}-\gamma Y^{\prime}_{i}X^{\prime}_{i}+\left(I-\gamma X^{\prime}_{i}X_{i}^{\prime\top}\right)\beta_{i-1}$
@@
$\eta_{i}$ $=(I-\gamma\|X_{i}\|^{2}H_{i})\eta_{i-1}$
$=H_{i}^{\perp}\eta_{i-1}+(1-\gamma\|X_{i}\|^{2})H_{i}\eta_{i-1},$
@@
$\|\eta_{i}\|^{2}$ $=\|H_{i}^{\perp}\eta_{i-1}\|^{2}+(1-\gamma\|X_{i}\|^{2})^{2}\|H_{i}\eta_{i-1}\|^{2}$
$\stackrel{{\scriptstyle(I)}}{{\leq}}\|H_{i}^{\perp}\eta_{i-1}\|^{2}+(1-\gamma\|X_{i}\|^{2})\|H_{i}\eta_{i-1}\|^{2}$
$\leq\|\eta_{i-1}\|^{2}-\gamma\|X_{i}\|^{2}\|H_{i}\eta_{i-1}\|^{2}$
@@
$\frac{\|\eta_{i}\|^{2}}{\|\eta_{i-1}\|^{2}}\leq 1-\gamma\|X_{i}\|^{2}\frac{\|H_{i}\eta_{i-1}\|^{2}}{\|\eta_{i-1}\|^{2}}=1-\gamma\frac{\eta_{i-1}^{\top}X_{i}X_{i}^{\top}\eta_{i-1}}{\|\eta_{i-1}\|^{2}}$
@@
$\mathbb{E}\left[\left.\frac{\|\eta_{i}\|^{2}}{\|\eta_{i-1}\|^{2}}\right|F^{i-1}\right]$ $\leq 1-\gamma\|\eta_{i-1}\|^{-2}\eta_{i-1}^{\top}\mathbb{E}[X_{i}X_{i}^{\top}]\eta_{i-1}$
@@
In step $(II)$ we used $\lambda_{min}(\mathbb{E}[X_{i}X_{i}^{\top}])\geq\underline{\lambda}>0$. Now we have established some “exponential contraction” properties for the sequence of vectors $\{\eta_{i}\}$. Then we can apply Lemma G.1 (with parameter $a$ therein equals to 00) to conclude that the averaged coefficient vector $\hat{\eta}_{i}=i^{-1}\sum_{k=1}^{i}\eta_{k}$ has magnitude:
$\mathbb{E}[\|\hat{\eta}_{i}\|^{2}|F^{j}]\lesssim i^{-2}\|\eta_{j}\|^{2}.$ (47)
@@
$\mathbb{E}\left[\left\{\nabla_{j}\hat{f}_{i-1}\left(X_{i}\right)\right\}^{2}\mid F^{j}\right]$ $=\mathbb{E}\left[\left\{\hat{f}_{i-1}\left(X_{i};Z_{i}\right)-\hat{f}_{i-1}\left(X_{i};Z_{i}^{j}\right)\right\}^{2}\mid F^{j}\right]$
$=\mathbb{E}\left[\left\{X_{i}^{\top}\hat{\beta}_{i-1}-X_{i}^{\top}\hat{\beta}_{i-1}^{\prime}\right\}^{2}\mid F^{j}\right]$
@@
$=\mathbb{E}\left[\left.\{X_{i}^{\top}\hat{\eta}_{i-1}\}^{2}\right|F^{j}\right]$
$\leq\mathbb{E}\left[\left.\|X_{i}\|^{2}\|\hat{\eta}_{i-1}\|^{2}\right|F^{j}\right],$
Under the assumption that $\mathbb{E}[\|X_{i}\|^{2}]\leq R^{2}$,
$\mathbb{E}\left[\left\{\nabla_{j}\hat{f}_{i-1}\left(X_{i}\right)\right\}^{2}\mid F^{j}\right]$ $\leq R^{2}\mathbb{E}\left[\left\|\hat{\eta}_{i-1}\right\|^{2}\mid F^{j}\right]$ (48)
@@
$\mathbb{E}\left[\left\{\nabla_{j}\hat{f}_{i-1}\left(X_{i}\right)\right\}^{2}\mid F^{j}\right]\lesssim i^{-2},$
@@
$\beta_{j}=\beta_{j-1}^{\wedge}+\gamma_{j}\left(Y_{j}-\bm{\phi}_{j}^{\top}\beta_{j-1}^{\wedge}\right)D_{j}\phi_{j},$
And we have a similar update rule for $\beta_{j}^{\prime}$ using sample $\left(X_{j}^{\prime},Y_{j}^{\prime}\right)$ :
$\beta_{j}^{\prime}$ $=\beta_{j-1}^{\prime\wedge}+\gamma_{j}\left(Y_{j}-\bm{\phi}_{j}^{\prime\top}\beta_{j-1}^{\prime\wedge}\right)D_{j}\bm{\phi}_{j}^{\prime}$
@@
$\eta_{j}=\gamma_{j}\left(Y_{j}-\bm{\phi}_{j}^{\top}\beta_{j-1}^{\wedge}\right)D_{j}\bm{\phi}_{j}-\gamma_{j}\left(Y_{j}-\bm{\phi}_{j}^{\prime\top}\beta_{j-1}^{\prime\wedge}\right)D_{j}\bm{\phi}_{j}^{\prime}.$
@@
$\eta_{i}=\eta_{i-1}^{\wedge}-\gamma_{i}D_{i}\bm{\phi}_{i}\bm{\phi}_{i}^{\top}\eta_{i-1}^{\wedge}$
@@
$D_{i}^{-1/2}\eta_{i}=D_{i}^{-1/2}\eta_{i-1}^{\wedge}-\gamma_{i}D_{i}^{1/2}\bm{\phi}_{i}\left(D_{i}^{1/2}\bm{\phi}_{i}\right)^{\top}D_{i}^{-1/2}\eta_{i-1}^{\wedge}$
@@
$\|D_{j}^{1/2}\bm{\phi}_{j}\|^{2}=\sum_{k=1}^{J_{j}}(k^{-\omega}\phi_{k}(X_{j}))^{2}\leq C\sum_{k=1}^{\infty}k^{-2\omega}<\infty,$
@@
$\left\|\theta_{i}\right\|^{2}$ $=\left\|\theta_{i-1}^{\wedge}\right\|^{2}+\gamma_{i}^{2}\left(\theta_{i-1}^{\wedge}\right)^{\top}W_{i}W_{i}^{\top}W_{i}W_{i}^{\top}\theta_{i-1}^{\wedge}-2\gamma_{i}\left(\theta_{i-1}^{\wedge}\right)^{\top}W_{i}W_{i}^{\top}\theta_{i-1}^{\wedge}$
@@
In step $(II)$ above we used $\gamma_{i}\|W_{i}\|^{2}\leq 1$. Take the conditional expectation
$\mathbb{E}[\|\theta_{i}\|^{2}|F^{i-1}]$ $\leq\|\theta^{\wedge}_{i-1}\|^{2}-\gamma_{i}\theta_{i-1}^{{}^{\wedge}\top}\mathbb{E}[W_{i}W_{i}^{\top}]\theta^{\wedge}_{i-1}$
$=\|\theta_{i-1}\|^{2}-\gamma_{i}\theta_{i-1}^{{}^{\wedge}\top}D_{i}^{1/2}\mathbb{E}[\bm{\phi}_{i}\bm{\phi}_{i}^{\top}]D_{i}^{1/2}\theta^{\wedge}_{i-1}$
@@
$\mathbb{E}[\|\theta_{i}\|^{2}/\|\theta_{i-1}\|^{2}|F^{i-1}]\leq 1-\gamma_{i}\frac{\theta_{i-1}^{{}^{\wedge}\top}D_{i}^{1/2}\mathbb{E}[\bm{\phi}_{i}\bm{\phi}_{i}^{\top}]D_{i}^{1/2}\theta^{\wedge}_{i-1}}{\|\theta^{\wedge}_{i-1}\|^{2}}$
By our assumption, $\lambda_{min}(\mathbb{E}[\bm{\phi}_{i}\bm{\phi}_{i}^{\top}])$ is strictly larger than a constant (we postpone the proof after equation (52)). Also note that $D_{i}^{1/2}={\rm Diag}(1^{-\omega},...,J_{i}^{-\omega})$ has eigenvalues greater than $J_{i}^{-\omega}=i^{-\zeta\omega}$. So overall we have
$\frac{\theta_{i-1}^{{}^{\wedge}\top}D_{i}^{1/2}\mathbb{E}[\bm{\phi}_{i}\bm{\phi}_{i}^{\top}]D_{i}^{1/2}\theta^{\wedge}_{i-1}}{\|\theta^{\wedge}_{i-1}\|^{2}}\geq i^{-2\zeta\omega}.$
Then we have an exponential contraction recursion:
$\mathbb{E}[\|\theta_{i}\|^{2}/\|\theta_{i-1}\|^{2}|F^{i-1}]\leq 1-\gamma_{i}i^{-2\zeta\omega}=1-i^{-2\zeta\omega-\zeta}\leq\exp(-i^{-\alpha}),$
@@
$\mathbb{E}[\|\hat{\theta}_{i}\|^{2}|F^{j}]\lesssim i^{-2}j^{2\alpha}\|\theta_{j}\|^{2}.$
@@
$\mathbb{E}[\|\hat{\theta}_{i}\|^{2}|F^{j}]\lesssim i^{-2}j^{2\alpha-2\zeta}=i^{-2}j^{4\omega\zeta}$